# Generalizing matrix representations to fully heterochronous ranked tree shapes

### v3

@@ 4.4 Simulations
For each simulated tree, we computed 3 statistics: the number of cherries, the total tree length (the sum of the number of $(u_{i},u_{i+1})$ intervals that each branch survives), and the internal tree length (the sum of the number of $(u_{i},u_{i+1})$ intervals that each internal edge survives). Since our models are models on tree topology only, we assumed a unit length interval between consecutive events (branching or sampling). The means of those statistics are presented in Tables 2 and 3. Empirical distributions based on 1000 simulations of trees with 20 leaves are depicted in Figure 6.
@@
Results from Tables 2 and 3 show that among the two parameter-free models, the coalescent model generates samples with larger average internal length, total length, and number of cherries compared to the diagonal top-down model. On the other hand, by adjusting the hyperparameters of the beta distribution in the Bernoulli splitting model, the resulting sample can be quite different in terms of the three average statistics.

### v4

@@ 4.4 Simulations
For each simulated tree, we computed 3 statistics: the number of cherries, the total tree length (the sum of the number of $(u_{i},u_{i+1})$ intervals that each branch survives), and the internal tree length (the sum of the number of $(u_{i},u_{i+1})$ intervals that each internal edge survives). Since our models are models on tree topology only, we assumed a unit length interval between consecutive events (branching or sampling). The means of those statistics are presented in Tables 3 and 4. Empirical distributions based on 1000 simulations of trees with 20 leaves are depicted in Figure 8.
@@
Results from Tables 3 and 4 show that among the two parameter-free models, the coalescent model generates samples with larger average internal length, total length, and number of cherries compared to the diagonal top-down model. On the other hand, by adjusting the hyperparameters of the beta distribution in the Bernoulli splitting model, the resulting sample can be quite different in terms of the three average statistics.